A tier 1 auto supplier is considering a new piece of equipment which costs $75,000. It is estimated to have a life of 15 years with a salvage value of $15,000 at that time. General operating expenses are expected to be $5,000/year, and repairs to the equipment are expected to be $750 the first year and increase by $150/year over the life of the equipment. An ROI of 12% is desired. Calculate the annual total cost of the equipment.
Answer #1

Here we need to calculate equivalent uniform annual cost (EUAC)

i = 12%

initial cost = 75000

Salvage value = 15000

life = 15 yrs

operating expenses = 5000/yr

repairs = 750 in first year increasing by 150 every year

EUAC = 75000*(A/P,12%,15) + 5000 + 750 + 150 *(A/G,12%,15) -15000*(A/F,12%,15)

= 75000*0.146824 + 5750 + 150 *4.980303 -15000*0.0268242

= 17106.50